OpenStudy (anonymous):

12x^2+4x-1 ------------ 2x+1

OpenStudy (anonymous):

Set something up like this: (2x+1) ( ) What goes on the far right to get 1*x = -1?

OpenStudy (anonymous):

That gives you (2x+1) ( - 1) 2x * -1 + (1 * Nx) = 4x. WHat is n?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

6x(2x+1)(2-1)?

OpenStudy (anonymous):

Here set it up like this: (2x+1)(Nx + y) The constant term of the result is -1, and you get the constant part by multiplying the last term of each binomial, so that gives: 1*y = -1 You solve for y to get -1 Now write (2x+1)(Nx + y) The x^2 part of the result is 12x^2, and you get the x^2 part by multiplying the first term of either binomial, so that gives: 2x*Nx = 12x^2 Solve for N. Or, like I was showing before, you get the x part of the result by: 2x*-1 + 1*Nx = 4x. Solve for N. Doing either of the last 2 should give you the right value for N.

OpenStudy (anonymous):

6x-1 is the answer right? the 2x+1 cancels out and 6x-1 is left ?